OpenStudy (condor34):

Can someone help with and Algebra 2 question? Solve 3(2x) = 7(x−1). Answer Choices: −7.74293 7.74293 −1 1

OpenStudy (chrisdbest):

A looks good

OpenStudy (condor34):

positive?

OpenStudy (chrisdbest):

For B: 46.45758 = 47.20051

OpenStudy (chrisdbest):

For A -14.74293 = -46.45758

OpenStudy (chrisdbest):

So Not A Or B

OpenStudy (chrisdbest):

For C: -6 = -8

OpenStudy (chrisdbest):

I just hope it's D... Or you worded it wrong.

OpenStudy (condor34):

im going to check if i worded it wrong brb

OpenStudy (chrisdbest):

For D: 6 = 0

OpenStudy (condor34):

Ya my bad 3^(2x) = 7^(x−1).

OpenStudy (chrisdbest):

L....O...to the mothafreakin L!

OpenStudy (chrisdbest):

So it is A!

OpenStudy (condor34):

thx